(a) Given a 100 MHz machine with a with a miss penalty of 20 cycles, a hit time of 2 cycles, and a miss rate of 5%, calculate the average memory access time (AMAT).
(b) Suppose doubling the size of the cache decrease the miss rate to 3%, but causes the hit time to increases to 3 cycles and the miss penalty to increase to 21 cycles. What is the
AMAT of the new machine?
Answer #1

a) Average memory-access time (AMAT) = Hit time + Miss rate x Miss penalty (ns or clocks)

AMAT = 2 + 0.05 * 20 = 3 clock cycles

b) Average memory-access time (AMAT) = Hit time + Miss rate x Miss penalty

AMAT = 3 + 0.03 * 21 = 3.63 clock cycles